OpenStudy (anonymous):

consider the following quadratic equation: 3x^2-13x-5=5 using the standard form ax^2+bx+c=0, of the equation above, factor the left hand side of the equation into two linear factors.

myininaya (myininaya):

3x^2-15x+2x-10=0 3x(x-5)+2(x-5)=0 (x-5)(3x+2)=0